Solve by using the quadratic formula (Put in standard form first)
2x2+3x=-7
the first part is 2x to the power of 2

Answer by mananth(16949)

2x2+3x=-7
2x^2+3x+7=0
compare with ax^2+bx+c
a=2,b=3,c=7
..
b^2-4ac=9-56=-47
x1= -3+sqrt(-47)/4
x1=-3/4 +sqrt(-47)/4
x1=-0.75+sqrt(-47)/4
x2=-0.75-sqrt(-47)/4